# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 2.5. Penny Forfeit
We now solve the one-step sub-game of $\mathrm{TLP}(\kappa,\rho)$, which we call $(\kappa,\rho)$-Penny Forfeit or $\mathrm{PF}(\kappa,\rho)$. In doing so, we will see the point of entry of the stronger condition $\rho\leq 1$, which found in the Nash-ABMN relationship as stated in Theorem 1.8.
@@
$\Big(\tfrac{\kappa a^{\rho}}{a^{\rho}+b^{\rho}}+\tfrac{1-\kappa}{2}\Big)m_{1}+\Big(\tfrac{\kappa b^{\rho}}{a^{\rho}+b^{\rho}}+\tfrac{1-\kappa}{2}\Big)m_{-1}-a\,\,\,\,\textrm{and}\,\,\,\,\Big(\tfrac{\kappa b^{\rho}}{a^{\rho}+b^{\rho}}+\tfrac{1-\kappa}{2}\Big)n_{-1}+\Big(\tfrac{\kappa a^{\rho}}{a^{\rho}+b^{\rho}}+\tfrac{1-\kappa}{2}\Big)n_{1}-b\,.$ (19)
@@
Suppose that $\rho\in(0,1]$. For $\kappa\in(0,1]$, there is a unique pair $(a,b)\in[0,\infty)^{2}$ for which the expressions in (19) are both global maxima as the variables $a$ and $b$ are respectively varied over $0,\infty)$. It is given by
$(a,b)\,=\,\kappa\rho\cdot\bigg(\frac{M^{1+\rho}N^{\rho}}{(M^{\rho}+N^{\rho})^{2}},\frac{M^{\rho}N^{1+\rho}}{(M^{\rho}+N^{\rho})^{2}}\bigg)\,,\,\,\,\,\textrm{with}\,\,\,\,M=m_{-1,1}\,\,\,\,\textrm{and}\,\,\,\,N=n_{1,-1}\,.$ (20)
@@
A critical point $(a,b)$ is given by setting the respective partial derivatives in $a$ and $b$ of the two expressions in (19) equal to zero: the conditions are
$\frac{\kappa\rho\,b^{\rho}a^{\rho-1}}{(a^{\rho}+b^{\rho})^{2}}M-1\,=\,\frac{\kappa\rho\,a^{\rho}b^{\rho-1}}{(a^{\rho}+b^{\rho})^{2}}N-1\,=\,0$ (21)
and these imply that $\kappa\rho\,b^{\rho}a^{\rho-1}M\;=\;\kappa\rho\,a^{\rho}b^{\rho-1}N$. Since $ab\neq 0$, $bM=aN$. Substituting $b=aN/M$ into $\kappa\rho\,b^{\rho}a^{\rho-1}M=(a^{\rho}+b^{\rho})^{2}$, dividing by $a^{2\rho-1}$ and rearranging yields the formula for $a$ in (20), with the formula for $b$ following from $b=aN/M$. The solution is positive and unique.
That the solution is a global maximum is due to $\rho\leq 1$. Indeed, $\frac{a^{\rho-1}}{(a^{\rho}+b^{\rho})^{2}}$ then has numerator that is decreasing in $a\geq 0$, so that, since the denominator is increasing in this variable, the expression is decreasing. With an analogous property for $\frac{b^{\rho-1}}{(a^{\rho}+b^{\rho})^{2}}$, this has the implication that the critical point in (21) is global in the sense of Lemma 2.10, completing the proof of this result. ∎
Remark. When $\rho>1$, the argument above continues to identify the pair $(a,b)$ in (21) as a critical point. However, the numerator in $\frac{a^{\rho-1}}{(a^{\rho}+b^{\rho})^{2}}$ is now increasing, and this sets up $a=0$ as a rival for the global maximizer of the first function in (19). The condition $M/N\geq(\rho-1)^{1/\rho}$ characterises when the rival $a=0$ falls short and when the putative critical point is global. Switching $M$ and $N$ in the last bound yields the applicable condition in regard to the second function in (19). In summary, a global maximum in the sense of Lemma 2.10 never exists when $\rho>2$; when $\rho\in(1,2]$, it entails that $M/N$ be suitably close to one, by lying in $\big[(\rho-1)^{1/\rho},(\rho-1)^{-1/\rho}\big]$.
@@
3. ABMN solutions: explicit forms and fixed-parameter asymptotics
Here we solve the ABMN($\kappa,\rho$) equations explicitly and deduce consequences. After giving the straightforward proof of the strict monotonicity of $m$- and $n$-differences recorded in Theorem 1.9(1), we re-express in the first subsection the ABMN system via the two-variable-per-site MN equations. This system permits the iterative computation of consecutive $m$- and $n$-differences, leading to the explicit sum-of-products representation in Theorem 1.16. The fixed-parameter asymptotics Theorem 1.21 will be obtained by analysing this representation. In the next two subsections, we offer elements needed for that analysis: first, the asymptotics of the map $s$; and then the resulting asmyptotics for the $s$-orbit. Obtaining also needed asymptotics for the $c$ and $d$ maps that appear in the products in the representation, we give the proof of Theorem 1.21 in the fourth subsection. The section ends with the proof of Theorem 1.9(2) on the finiteness of boundary data for elements of ABMN($\kappa,\rho$), which is a quick corollary of Theorem 1.21.
@@ 3.1. Explicit ABMN solutions
$a_{i}=\frac{\rho\kappa M_{i}^{1+\rho}N_{i}^{\rho}}{(M_{i}^{\rho}+N_{i}^{\rho})^{2}}\,\,\,\,,\,\,\,\,b_{i}=\frac{\rho\kappa M_{i}^{\rho}N_{i}^{1+\rho}}{(M_{i}^{\rho}+N_{i}^{\rho})^{2}}\,\,\,\,\textrm{and}\,\,\,\,\frac{a_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}=\frac{M_{i}^{\rho}}{M_{i}^{\rho}+N_{i}^{\rho}}\,.$ (22)
@@
Proof. From ABMN(3,4)(3,4) follows (22). Expressing ABMN(1)(1) in the form (33), we find from (22) that
@@
From Definition 1.12, recall the four basic functions $\gamma$, $\delta$, $\phi_{0}$ and $\phi_{1}$ that map $(\kappa,\rho,\beta)\in(0,1]^{2}\times(0,\infty)$ to $(0,\infty)$.
@@
In MN($2$), write $n_{i+1,i}=M_{i}\beta_{i}(1-\delta_{i})$, divide by $M_{i}^{2\rho+1}$, use $\beta_{i}=N_{i}/M_{i}$, cancel a factor of $\beta_{i}$ and rearrange, to obtain
@@
$\delta_{i}$ $=$ $\frac{(1-\kappa)\beta_{i}^{2\rho}+2(1+\rho\kappa)\beta_{i}^{\rho}+1+\kappa}{2(1+\beta_{i}^{\rho})^{2}}\,.$
@@
$c_{i}(x)-1\,=\,\frac{1-\gamma(\kappa,\rho,\beta_{i})}{\gamma(\kappa,\rho,\beta_{i})}\,=\,\frac{1-\gamma_{i}}{\gamma_{i}}\,=\,\frac{m_{i,i+1}}{m_{i,i-1}}\,$, (23)
the second equality [^2] Usages such as Lemma 3.3($\gamma$) and Proposition 5.4(2,$f$) refer to the statement made the result in question about the object $\gamma$ or $f$. due to Lemma 3.3($\gamma$) and the third to Definition 3.2($\gamma$). With the product notation from Definition 1.15 applying negative index $j$, we find that
$\frac{m_{j,j+1}}{m_{-1,0}}\,=\,\prod_{i=0}^{j}\big(c_{i}(x)-1\big)$ (24)
for any $j\in\mathbb{Z}$. The ABMN($\kappa,\rho$) element $(a,b,m,n)$ under consideration may be dilated by varying $m_{-1,0}\in(0,\infty)$, in correspondence with the dilation factor $\mu$ that appears in Theorem 1.16. By setting $m_{-1,0}=1$, we reduce the task of proving the theorem to checking that $(a,b,m,n)$ equals the default quadruple $\big(a^{\rm def}(x),b^{\rm def}(x),m^{\rm def}(x),n^{\rm def}(x)\big)$ (so $\mu=1$). And indeed we have proved the $m$-component projection of the desired identity, because the right-hand side in (24) is $m^{\rm def}_{k+1}(x)-m^{\rm def}_{k}(x)$ from Definition 1.15.
Evident variations of the argument leading to (23) yield
@@ 3.2. $\phi_{0}$, $\phi_{1}$ and $s$ asymptotics
Proof. In either case, the weaker condition we consider for a $(\kappa,\rho)$ pair, namely membership of $W$ as specified in (8), is met; this enables the use of Lemma 2.3, so $s$ is well defined. The $\phi$-asymptotics are computed by working with the formulas in Definition 1.12. For example, in the latter case,
$\phi_{0}(\beta)$ $=\beta\,\frac{2\beta^{2\rho}\big(1+\tfrac{2\beta^{\rho}-(1-\rho)}{2\beta^{2\rho}}\big)}{2\beta^{\rho}\big(1+\tfrac{1-\rho}{\beta^{\rho}}\big)}=\beta\,\frac{\beta^{2\rho}}{\beta^{\rho}}\cdot\frac{1+O(\beta^{-\rho})}{1+O(\beta^{-\rho})}=\frac{1+\rho}{1-\rho}\,\beta+O(\beta^{1-\rho}),$
$\phi_{1}(\beta)$ $=\beta\,\frac{\beta^{\rho}(1+(1-\rho)/\beta^{\rho})}{\beta^{\rho}(1+(1+\rho)/\beta^{\rho})}=\beta\,\frac{1+O(\beta^{-\rho})}{1+O(\beta^{-\rho})}=\beta+O(\beta^{1-\rho}).$
@@ 3.3. Asymptotics for the backward orbit of $s$
$s_{-1}(x)=x+\frac{8\rho\kappa}{1-\kappa^{2}}\,x^{1-\rho}+O\left(x^{1-2\rho}\right)\quad\text{as }x\to\infty,$ (25)
@@
We give a rigorous argument for $\rho\in(0,1)$; the $\rho=1$ involves introducing suitable logarithmic factors. Find large positive constants $C_{0}$ and $D$ such that for $x\geq C_{0}$ the implied constant in the big-$O$ term in (25) is at most $D$. By orbit escape Corollary 2.6, select $n_{0}$ such that $\phi_{-n}\geq C_{0}$ for $n\geq n_{0}$. For any $n$, there exists an integer $l=l(n)$ such that
$\big|\phi_{-n}-A(n+l)^{1/\rho}\big|\leq 2A\rho^{-1}(n+l)^{\rho^{-1}-1}$
where recall that $A$ equals $(C\rho)^{1/\rho}$. (The right-hand factor of two copes with increase associated to making $l(n)$ an integer.) For a positive constant $K$ suitably determined by $D$, we select $n_{1}\geq\max\{n_{0},K^{2}\}$, and set $\ell=l(n_{1})$ (so that the offset value $\ell\in\mathbb{N}$ is now fixed, independently of $n$). Setting $x_{n}=\phi_{-(n-\ell)}$ and $e_{n}=x_{n}-An^{1/\rho}$, we will argue by induction on $n\geq n_{1}+\ell$ that $\big|e_{n}\big|\leq K\,n^{\rho^{-1}-1}$. Expanding the power of $i+\ell$ in the resulting upper bound on $|e_{i+\ell}|$ yields the claim asymptotic on $\phi_{-i}$.
@@
where the displayed equality is due the inductive hypothesis in the guise $n^{-1/\rho-1}e_{n}^{2}\leq K^{2}n^{\rho^{-1}-3}$ and $n^{-\rho^{-1}}e_{n}\leq Kn^{-1}$. The final displayed term may be written $O(1)n^{\rho^{-1}-2}$ since $n\geq n_{1}+\ell\geq n_{1}\geq K^{2}$. Substituting back into (26), and noting that the resulting right-hand $CA^{1-\rho}n^{\rho^{-1}-1}$ term equals
@@
In regard to $\beta$-asymptotics, note that $\beta_{i}=\beta(\phi_{i})$, so that $\phi_{i}=\beta_{i}\big(1+O(\beta_{i}^{-\rho})\big)$ whenever $\rho\in(0,1]$, by Lemma 3.4(1); consequently $\beta_{i}=\phi_{i}\big(1+O(\phi_{i}^{-\rho})\big)$ . So the $\phi$-asymptotics pass to the $\beta$-sequence.

### v4

@@ 2.5. Penny Forfeit
We now solve the one-step sub-game of $\mathrm{TLP}(\kappa,\rho)$, which we call $(\kappa,\rho)$-Penny Forfeit or $\mathrm{PF}(\kappa,\rho)$. In doing so, we will see the point of entry of the stronger condition $\rho\leq 1$, which is found in the Nash-ABMN relationship as stated in Theorem 1.3.
@@
$\Big(\tfrac{\kappa a^{\rho}}{a^{\rho}+b^{\rho}}+\tfrac{1-\kappa}{2}\Big)m_{1}+\Big(\tfrac{\kappa b^{\rho}}{a^{\rho}+b^{\rho}}+\tfrac{1-\kappa}{2}\Big)m_{-1}-a\,\,\,\,\textrm{and}\,\,\,\,\Big(\tfrac{\kappa b^{\rho}}{a^{\rho}+b^{\rho}}+\tfrac{1-\kappa}{2}\Big)n_{-1}+\Big(\tfrac{\kappa a^{\rho}}{a^{\rho}+b^{\rho}}+\tfrac{1-\kappa}{2}\Big)n_{1}-b\,.$ (2.2)
@@
Suppose that $\rho\in(0,1]$. For $\kappa\in(0,1]$, there is a unique pair $(a,b)\in[0,\infty)^{2}$ for which the expressions in (2.2) are both global maxima as the variables $a$ and $b$ are respectively varied over $0,\infty)$. It is given by
$(a,b)\,=\,\kappa\rho\cdot\bigg(\frac{M^{1+\rho}N^{\rho}}{(M^{\rho}+N^{\rho})^{2}},\frac{M^{\rho}N^{1+\rho}}{(M^{\rho}+N^{\rho})^{2}}\bigg)\,,\,\,\,\,\textrm{with}\,\,\,\,M=m_{-1,1}\,\,\,\,\textrm{and}\,\,\,\,N=n_{1,-1}\,.$ (2.3)
@@
A critical point $(a,b)$ is given by setting the respective partial derivatives in $a$ and $b$ of the two expressions in (2.2) equal to zero: the conditions are
$\frac{\kappa\rho\,b^{\rho}a^{\rho-1}}{(a^{\rho}+b^{\rho})^{2}}M-1\,=\,\frac{\kappa\rho\,a^{\rho}b^{\rho-1}}{(a^{\rho}+b^{\rho})^{2}}N-1\,=\,0$ (2.4)
and these imply that $\kappa\rho\,b^{\rho}a^{\rho-1}M\;=\;\kappa\rho\,a^{\rho}b^{\rho-1}N$. Since $ab\neq 0$, $bM=aN$. Substituting $b=aN/M$ into $\kappa\rho\,b^{\rho}a^{\rho-1}M=(a^{\rho}+b^{\rho})^{2}$, dividing by $a^{2\rho-1}$ and rearranging yields the formula for $a$ in (2.3), with the formula for $b$ following from $b=aN/M$. The solution is positive and unique.
That the solution is a global maximum is due to $\rho\leq 1$. Indeed, $\frac{a^{\rho-1}}{(a^{\rho}+b^{\rho})^{2}}$ then has numerator that is decreasing in $a\geq 0$, so that, since the denominator is increasing in this variable, the expression is decreasing. With an analogous property for $\frac{b^{\rho-1}}{(a^{\rho}+b^{\rho})^{2}}$, this has the implication that the critical point in (2.4) is global in the sense of Lemma 2.10, completing the proof of this result. ∎
Other parameter regimes
When $\rho>1$, the argument above continues to identify the pair $(a,b)$ in (2.4) as a critical point. However, the numerator in $\frac{a^{\rho-1}}{(a^{\rho}+b^{\rho})^{2}}$ is now increasing, and this sets up $a=0$ as a rival for the global maximizer of the first function in (2.2). The condition $M/N\geq(\rho-1)^{1/\rho}$ characterises when the rival $a=0$ falls short and when the putative critical point is global. Switching $M$ and $N$ in the last bound yields the applicable condition in regard to the second function in (2.2). In summary, a global maximum in the sense of Lemma 2.10 never exists when $\rho>2$; when $\rho\in(1,2]$, it entails that $M/N$ be suitably close to one, by lying in $\big[(\rho-1)^{1/\rho},(\rho-1)^{-1/\rho}\big]$.
@@
‘Hand It Over!’: A strategically equivalent zero-sum game when terminal rewards are equal
Moulin and Vial [40] developed a theory reducing certain two-person games to strategically equivalent zero-sum games, in the sense that equilibrium strategies coincide up to constant payoff shifts. This framework is applicable to ${\rm PF}(\kappa,\rho)$ in the equal-reward case $m_{-1,1}=n_{1,-1}$.
In the strategically equivalent variant ‘Hand It Over!’ game, stakes are not surrendered to the bank; rather, the stake of a given player is handed to the opponent. After suitable constant payments are added to each player’s payoff, a zero-sum game results whose equilibria coincide with those of ${\rm PF}(\kappa,\rho)$. Indeed, the respective mean winnings in (2.2) are shifted by amounts that are independent of $a$ and of $b$, leaving the solutions in (2.3) unaltered.
This strategic equivalence breaks as soon as the gameboard has at least two sites in open play, because the equilibrium values $m_{i}$ and $n_{i}$ then differ from those of the one-step game by terms that depend nontrivially on equilibrium stakes from both players.
Chapter 3 ABMN solutions: explicit forms and fixed-parameter asymptotics
Here we solve the ABMN($\kappa,\rho$) equations explicitly and deduce consequences. After giving the straightforward proof of the strict monotonicity of $m$- and $n$-differences recorded in Theorem 1.4(1), we re-express in the first section the ABMN system via the two-variable-per-site MN equations. This system permits the iterative computation of consecutive $m$- and $n$-differences, leading to the explicit sum-of-products representation in Theorem 1.11. The fixed-parameter asymptotics Theorem 1.22 will be obtained by analysing this representation. In the next two sections, we offer elements needed for that analysis: first, the asymptotics of the map $s$; and then the resulting asymptotics for the $s$-orbit. Obtaining also needed asymptotics for the $c$ and $d$ maps that appear in the products in the representation, we give the proof of Theorem 1.22 in the fourth section. The chapter ends with the proof of Theorem 1.4(2) on the finiteness of boundary data for elements of ABMN($\kappa,\rho$), which is a quick corollary of Theorem 1.22.
@@ 3.1. Explicit ABMN solutions
$a_{i}=\frac{\rho\kappa M_{i}^{1+\rho}N_{i}^{\rho}}{(M_{i}^{\rho}+N_{i}^{\rho})^{2}}\,\,\,\,,\,\,\,\,b_{i}=\frac{\rho\kappa M_{i}^{\rho}N_{i}^{1+\rho}}{(M_{i}^{\rho}+N_{i}^{\rho})^{2}}\,\,\,\,\textrm{and}\,\,\,\,\frac{a_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}=\frac{M_{i}^{\rho}}{M_{i}^{\rho}+N_{i}^{\rho}}$ (3.1)
@@
Proof. From ABMN(3,4)(3,4) follows (3.1). Since $a_{i}^{\rho}+b_{i}^{\rho}>0$, we may express ABMN(1)(1) in the form
$m_{i}\,=\,\Big(\kappa\tfrac{a_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}+\tfrac{1-\kappa}{2}\Big)m_{i+1}+\Big(\kappa\tfrac{b_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}+\tfrac{1-\kappa}{2}\Big)m_{i-1}\,-\,a_{i}\,$,
we find from (3.1) that
@@
From Definition 1.7, recall the four basic functions $\gamma$, $\delta$, $\phi_{0}$ and $\phi_{1}$ that map $(\kappa,\rho,\beta)\in W\times(0,\infty)$ to $(0,\infty)$.
@@
In MN($2$), write $n_{i+1,i}=M_{i}\beta_{i}(1-\delta_{i})$, divide by $M_{i}^{2\rho+1}$, use $\beta_{i}=N_{i}/M_{i}$, cancel a factor of $\beta_{i}$ and rearrange, to arrive at
@@
$\delta_{i}$ $=$ $\frac{(1-\kappa)\beta_{i}^{2\rho}+2(1+\rho\kappa)\beta_{i}^{\rho}+1+\kappa}{2(1+\beta_{i}^{\rho})^{2}}\,$,
@@
$c_{i}(x)-1\,=\,\frac{1-\gamma(\kappa,\rho,\beta_{i})}{\gamma(\kappa,\rho,\beta_{i})}\,=\,\frac{1-\gamma_{i}}{\gamma_{i}}\,=\,\frac{m_{i,i+1}}{m_{i,i-1}}\,$, (3.2)
the second equality [^1] Usages such as Lemma 3.3($\gamma$) and Proposition 5.4(2,$f$) refer to the statement made the result in question about the object $\gamma$ or $f$. due to Lemma 3.3($\gamma$) and the third to Definition 3.2($\gamma$). With the product notation from Definition 1.10 applying when the index $j$ is negative, we find that
$\frac{m_{j,j+1}}{m_{-1,0}}\,=\,\prod_{i=0}^{j}\big(c_{i}(x)-1\big)$ (3.3)
for any $j\in\mathbb{Z}$. The ABMN($\kappa,\rho$) element $(a,b,m,n)$ under consideration may be dilated by varying $m_{-1,0}\in(0,\infty)$, in correspondence with the dilation factor $\mu$ that appears in Theorem 1.11. By setting $m_{-1,0}=1$, we reduce the task of proving the theorem to checking that $(a,b,m,n)$ equals the default quadruple $\big(a^{\rm def}(x),b^{\rm def}(x),m^{\rm def}(x),n^{\rm def}(x)\big)$ (so $\mu=1$). And indeed we have proved the $m$-component projection of the desired identity, because the right-hand side in (3.3) is $m^{\rm def}_{k+1}(x)-m^{\rm def}_{k}(x)$ from Definition 1.10.
Evident variations of the argument leading to (3.2) yield
@@ 3.2. $\phi_{0}$, $\phi_{1}$ and $s$ asymptotics
Proof. In either case, the weaker condition we consider for a $(\kappa,\rho)$ pair, namely membership of $W$ as specified in (1.7), is met; this enables the use of Lemma 2.1, so $s$ is well defined. The $\phi$-asymptotics are computed by working with the formulas in Definition 1.7. For example, in the latter case $\kappa=1$ and $\rho\in(0,1)$, the formulas (1.9) and (1.10) reduce to
$\phi_{0}(1,\rho,\beta)=\beta\cdot\frac{2(1+\rho)\beta^{\rho}+2}{2(1-\rho)\beta^{\rho}+2}=\beta\cdot\frac{(1+\rho)\beta^{\rho}+1}{(1-\rho)\beta^{\rho}+1}$
and
$\phi_{1}(1,\rho,\beta)=\beta\cdot\frac{2\beta^{2\rho}+2(1-\rho)\beta^{\rho}}{2\beta^{2\rho}+2(1+\rho)\beta^{\rho}}=\beta\cdot\frac{\beta^{\rho}+(1-\rho)}{\beta^{\rho}+(1+\rho)}\,.$
Thus, as $\beta\to\infty$,
$\phi_{0}(1,\rho,\beta)=\tfrac{1+\rho}{1-\rho}\,\beta+O(\beta^{1-\rho})\qquad\text{and}\qquad\phi_{1}(1,\rho,\beta)=\beta+O(\beta^{1-\rho})\,.$
@@ 3.3. Asymptotics for the backward orbit of $s$
$s_{-1}(x)=x+\frac{8\rho\kappa}{1-\kappa^{2}}\,x^{1-\rho}+O\left(x^{1-2\rho}\right)\quad\text{as }x\to\infty,$ (3.4)
@@
We give a rigorous argument for $\rho\in(0,1)$; the $\rho=1$ involves introducing suitable logarithmic factors. Find large positive constants $C_{0}$ and $D$ such that for $x\geq C_{0}$ the implied constant in the big-$O$ term in (3.4) is at most $D$. By orbit escape Corollary 2.4, select $n_{0}$ such that $\phi_{-n}\geq C_{0}$ for $n\geq n_{0}$. Set $t_{n}=(\phi_{-n}/A)^{\rho}-n$, so that $\phi_{-n}=A(n+t_{n})^{1/\rho}$. Let $l=l(n)\in\mathbb{Z}$ be the nearest integer to $t_{n}$. Since $\phi_{-n}\geq C_{0}$, for $n\geq n_{0}$, the unit neighbourhood of $n+t_{n}$ lies in $(0,\infty)$, so that we may apply the mean value theorem to the map $u\mapsto A(n+u)^{1/\rho}$ to learn that
$\big|\phi_{-n}-A(n+l(n))^{1/\rho}\big|=A\big|(n+t_{n})^{1/\rho}-(n+l(n))^{1/\rho}\big|\leq 2A\rho^{-1}(n+l)^{\rho^{-1}-1}$
where recall that $A$ equals $(C\rho)^{1/\rho}$. For a positive constant $K$ suitably determined by $D$, we select $n_{1}\geq\max\{n_{0},K^{2}\}$, and set $\ell=l(n_{1})$ (so that the offset value $\ell\in\mathbb{N}$ is now fixed, independently of $n$). Setting $x_{n}=\phi_{-(n-\ell)}$ and $e_{n}=x_{n}-An^{1/\rho}$, we will argue by induction on $n\geq n_{1}+\ell$ that $\big|e_{n}\big|\leq K\,n^{\rho^{-1}-1}$. Expanding the power of $i+\ell$ in the resulting upper bound on $|e_{i+\ell}|$ yields the claim asymptotic on $\phi_{-i}$.
@@
where the displayed equality is due the inductive hypothesis in the guise $n^{-1/\rho-1}e_{n}^{2}\leq K^{2}n^{\rho^{-1}-3}$ and $n^{-\rho^{-1}}e_{n}\leq Kn^{-1}$. The final displayed term may be written $O(1)n^{\rho^{-1}-2}$ since $n\geq n_{1}+\ell\geq n_{1}\geq K^{2}$. Substituting back into (3.5), and noting that the resulting right-hand $CA^{1-\rho}n^{\rho^{-1}-1}$ term equals
@@
In regard to $\beta$-asymptotics, note that $\beta_{i}=\beta(\phi_{i})$, so that $\phi_{i}=\beta_{i}\big(1+O(\beta_{i}^{-\rho})\big)$ whenever $\kappa\in(0,1)$ and $\rho\in(0,1]$, by Lemma 3.4(1); consequently $\beta_{i}=\phi_{i}\big(1+O(\phi_{i}^{-\rho})\big)$ . So the $\phi$-asymptotics pass to the $\beta$-sequence.